Determine the truth value for each of the following statements:
(i) 4 + 3 = 6 AND 3 + 3 = 6
(ii) 5 + 3 = 8 OR 3 + 1 = 5

Answer by jim_thompson5910(35256)

i) 4 + 3 = 6 AND 3 + 3 = 6 is false because 4+3=6 is false (both need to be true in order for entire statement to be true)

ii) 5 + 3 = 8 OR 3 + 1 = 5 is true since 5+3=8 is true (you need at least one statement to be true for entire statement to be true)


If you need more help, email me at jim_thompson5910@hotmail.com

Also, feel free to check out my tutoring website

Jim